# Joint Self-Supervised Video Alignment and Action Segmentation

### v2

@@ 3.1.3 Self-Supervised Learning
with $\tau>0$ denoting a temperature scaling parameter. In addition, pseudo-labels $\mathbf{T}^{\star}$ are obtained by solving the balanced FGW problem in Eq. 3 but with the augmented KOT cost matrix $\tilde{\mathbf{C}}=\mathbf{C}+\rho\mathbf{R}$, with $\rho\geq 0$. Here, we follow VAVA [45] to impose the temporal prior $\mathbf{R}$, defined as $\mathbf{R}_{ij}=\left|i/N-j/M\right|$, which encourages the coupling $\mathbf{T}$ to have a banded diagonal shape and temporally nearby frames in $X$ to be matched with temporally adjacent frames in $Y$, yielding improved performance as seen in Sec. 4.1. Finally, our self-supervised learning loss is written as:
@@ 3.2 Joint Self-Supervised Video Alignment and Action Segmentation
Sec. 3.1 presents our self-supervised video alignment approach (VAOT) developed based on an FGW optimal transport formulation with a structural prior, which has previously been adopted for self-supervised action segmentation by ASOT [80]. Below we discuss differences between self-supervised video alignment and action segmentation, which lead to distinct design choices for VAOT and ASOT [80]. Firstly, as illustrated in Fig. 1, video alignment performs finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation, which causes our removal of $\mathbf{A}$ and our cost matrices $\{\mathbf{C},\mathbf{C}^{x},\mathbf{C}^{y}\}$ in Sec. 3.1 to be different from ASOT [80]. Secondly, video alignment generally has a more balanced assignment than action segmentation (the number of frames is much larger than the number of actions) and it is hard to balance multiple losses (the full unbalanced formulation adds two extra penalty terms). Thus, a balanced FGW formulation performs the best for VAOT, whereas a partial unbalanced FGW formulation is preferred in ASOT [80]. Thirdly, a background action class, to which background/redundant frames are assigned, is typically included in the $K$ actions for action segmentation, whereas it is not already defined for video alignment. Thus, VAOT adds a virtual frame to tackle background/redundant frames.
@@ 3.2.2 Self-Supervised Multi-Task Learning
Since both self-supervised video alignment and action segmentation exploit fine-grained temporal information in videos, we propose a self-supervised multi-task learning framework for joint video alignment and action segmentation. In particular, we combine our VAOT module for video alignment with ASOT [80] for action segmentation into a unified optimal transport-based approach (VASOT), which is illustrated in Fig. 2(b). Here, we update variable names in VAOT from $\mathbf{C},\mathbf{P},\mathbf{T}^{\star},\mathcal{L}$} to $\mathbf{C}^{xy},\mathbf{P}^{xy},\mathbf{T}^{xy\star},\mathcal{L}_{xy}$} respectively for video alignment between $X$ and $Y$, while introducing new variables $\mathbf{C}^{xa},\mathbf{P}^{xa},\mathbf{T}^{xa\star},\mathcal{L}_{xa}$} and $\mathbf{C}^{ya},\mathbf{P}^{ya},\mathbf{T}^{ya\star},\mathcal{L}_{ya}$} for action segmentation on $X$ and $Y$ respectively. The parameters $\bm{\theta}$ of the frame encoder and the action embeddings $\mathbf{A}$ are trained by using the below combination of self-supervised learning losses:
@@
where $w_{align}\geq 0$ and $w_{seg}\geq 0$ denote the weights for the video alignment loss $\mathcal{L}_{xy}$ and the action segmentation losses $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively, $\mathcal{L}_{xy}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xy}$ (computed between $\mathbf{X}$ and $\mathbf{Y}$) and pseudo-labels $\mathbf{T}^{xy\star}$ (obtained by VAOT), while $\mathcal{L}_{xa}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xa}$ (computed based on $\mathbf{X}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{xa\star}$ (derived by ASOT [80]) and $\mathcal{L}_{ya}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{ya}$ (computed between $\mathbf{Y}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{ya\star}$ (obtained by ASOT [80]). This is in contrast with VAOT or ASOT [80], where ${\bm{\theta}}$ is solely learned by using either $\mathcal{L}_{xy}$ or $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively. We find balancing $w_{align}=w_{seg}=1$ yields good results for both video alignment and action segmentation, as seen in Sec. 4.2. Our joint model requires training and storing a single model, saving both time and memory usage as compared to two single-task models. As we observe in Sec. 4.3, in a multi-task learning setting, action segmentation offers little benefit to video alignment performance, while video alignment boosts action segmentation results substantially.
@@ 4 Experiments
Datasets. We benchmark our VAOT and VASOT approaches for video alignment using three datasets, including monotonic datasets, i.e., Pouring [57] and Penn Action [82], and in-the-wild dataset, i.e., IKEA ASM [3]. Pouring includes videos of humans pouring liquids and Penn Action comprises of videos of humans playing sports, while IKEA ASM videos capture humans assembling furniture. All methods have the same training and validation splits. Moreover, for action segmentation evaluation, we use four datasets, including in-the-wild datasets, i.e., Breakfast [33], 50 Salads [62], and YouTube Instructions [2], and monotonic dataset, i.e., Desktop Assembly [35]. Breakfast and 50 Salads videos show cooking activities, while YouTube Instructions consists of instructional videos and Desktop Assembly includes videos of an assembly activity. All methods are trained and tested on the same set of videos. For 50 Salads, we evaluate at two action granularity levels, i.e., Mid with 19 actions and Eval with 12 actions. Finally, for datasets with many activities, i.e., Penn Action, Breakfast, and YouTube Instructions, we train and test the methods per activity and report the average results.
@@
Implementation Details. For fair comparison purposes, our VAOT and VASOT approaches for video alignment utilize the same ResNet-50 encoder as recent self-supervised video alignment methods [19, 26, 45, 17]. Similarly, our VASOT approach for action segmentation employs the same MLP encoder as state-of-the-art self-supervised action segmentation methods [34, 35, 80]. Action embeddings $\mathbf{A}$ are initialized via $K$-Means, while the number of clusters $K$ is set to the ground truth value. We implement our methods in PyTorch [51] and use ADAM optimization [32]. Please refer to our supplementary material for more details.
@@
Competing Methods. We compare our VAOT and VASOT approaches against prior self-supervised video alignment methods, namely SAL [48], TCN [57], TCC [19], LAV [26], VAVA [45], and GTCC [17]. VAVA [45], which integrates an optimality prior into a classical Kantorovich optimal transport, is the closest to our VAOT approach. Also, we test our VASOT approach against previous self-supervised action segmentation methods, namely CTE [34], VTE [73], UDE [65], ASAL [43], TOT [35], UFSA [71], ASOT [80], and HVQ [60]. ASOT [80] is the single-task baseline for action segmentation, which we adopt in Sec. 3.2 for our multi-task VASOT approach.
@@
Evaluation Metrics. To evaluate our VAOT and VASOT approaches for video alignment, we compute four metrics on the validation set, i.e., phase classification (Acc@{0.1,0.5,1.0}), phase progression (Progress), video alignment ($\tau$), and fine-grained frame retrieval (AP@{5,10,15}). Prior to that, we train the model on the training set and freeze it, and then train an SVM classifier or linear regressor on top of frozen features. Note that as mentioned in [19, 26, 45, 17], Progress and $\tau$ are only defined for monotonic datasets and hence are not computed for IKEA ASM. Also, to test our VASOT approach for action segmentation, we calculate three metrics, i.e., mean over frames (MoF), F1 score (F1), and mean intersection over union (mIoU). Before that, we train the model, obtain predicted action segments, and perform Hungarian matching between predicted and ground truth action clusters.
@@ 4.2 Sensitivity Analysis Results
(e) Segmentation loss weight $w_{seg}$ (here, $w_{align}=1$).
Here, we conduct sensitivity analyses on hyperparameters of VAOT and VASOT. Fig. 3(d) shows the results. We use Pouring in Figs. 3(d)(a-e) and Desktop Assembly in Fig. 3(d)(f). The results of $\epsilon$ are provided in our supplementary material.
Effect of $r$ and $\alpha$. From Fig. 3(d)(a), Acc@1.0 remains stable and Progress shows small variation across all studied values of $r$, whereas $\tau$ is the most sensitive metric, peaking at $r=0.02$ and decreasing as $r$ increases. Similar observations can be made for $\alpha$ in Fig. 3(d)(b), where Acc@1.0 and Progress are mostly stable, whereas $\tau$ fluctuates the most, performing the best with $\alpha=0.3$. Furthermore, we find that $r=0.02$ and $\alpha=0.3$ also work the best for the remaining datasets.
Effect of $\rho$ and $\zeta$. For $\rho$ in Fig. 3(d)(c), it can be seen that Acc@1.0 and Progress remain mostly stable, whereas $\tau$ varies the most, performing the best at $\rho=0.35$. Similarly, for $\zeta$ in Fig. 3(d)(d), we observe that $\tau$ is the most sensitive metric, peaking at $\zeta=0.5$, whereas Acc@1.0 and Progress remain steady across the analyzed value range of $\zeta$. Moreover, we notice that $\rho=0.35$ and $\zeta=0.5$ also yield the best results for the remaining datasets.
Effect of $w_{seg}$ and $w_{align}$. We study the relationship between action segmentation and video alignment in our multi-task VASOT approach by varying their weights $w_{seg}$ and $w_{align}$ in Eq. 7. In particular, we report alignment results with various $w_{seg}$ values (while keeping $w_{align}=1$) in Fig. 3(d)(e), and segmentation results with various $w_{align}$ values (while keeping $w_{seg}=1$) in Fig. 3(d)(f). It can be seen from Fig. 3(d)(e) that alignment results drop as $w_{seg}$ increases and the best overall alignment performance is obtained with $w_{seg}=1$. In contrast, it is clear from Fig. 3(d)(f) that segmentation results improve as $w_{align}$ increases and become steady after $w_{align}$ reaches $1$. Therefore, in a multi-task learning setup, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance notably. Moreover, with balancing $w_{seg}=w_{align}=1$, VASOT achieves good results for both video alignment and action segmentation. For the next section, we set $w_{seg}=w_{align}=1$ for VASOT.
@@ 4.3 State-of-the-Art Comparison Results
Pouring SAL [48] 85.68 87.84 88.02 74.51 73.31 84.05 83.77 83.79
TCN [57] 89.19 90.39 90.35 80.57 86.69 83.56 83.31 83.01
TCC [19] 89.23 91.43 91.82 80.30 85.16 87.16 86.68 86.54
LAV [26] 91.61 92.82 92.84 80.54 85.61 89.13 89.13 89.22
VAVA [45] 91.65 91.79 92.45 83.61 87.55 90.05 89.92 90.17
GTCC [17] 71.20 89.20 93.50 85.80 88.10 - - -
@@
Penn Action SAL [48] 74.87 78.26 79.96 59.43 63.36 76.04 75.77 75.61
TCN [57] 81.99 83.67 84.04 67.62 73.28 77.84 77.51 77.28
TCC [19] 81.26 83.35 84.45 67.26 73.53 76.74 76.27 75.88
LAV [26] 83.56 83.95 84.25 66.13 80.47 79.13 78.98 78.90
VAVA [45] 83.89 84.23 84.48 70.91 80.53 81.52 80.47 80.67
GTCC [17] 78.30 81.20 81.30 70.80 88.30 - - -
@@
IKEA ASM SAL [48] 22.94 23.43 25.46 - - 14.28 14.04 14.10
TCN [57] 22.51 25.47 25.88 - - 17.37 17.03 16.96
TCC [19] 22.70 25.04 25.63 - - 18.03 17.53 17.20
LAV [26] 23.19 25.47 25.54 - - 20.14 19.35 19.21
VAVA [45] 29.12 29.95 29.10 - - 26.42 25.73 25.80
@@
Full-Dataset Evaluation CTE* [34] 41.8 / 26.4 / - 39.0 / 28.3 / - 30.2 / - / - 35.5 / - / - 47.6 / 44.9 / -
CTE† [34] 47.2 / 27.0 / 14.9 35.9 / 28.0 / 9.9 30.1 / 25.5 / 17.9 35.0 / 35.5 / 21.6 - / - / -
VTE [73] 48.1 / - / - - / 29.9 / - 24.2 / - / - 30.6 / - / - - / - / -
UDE [65] 47.4 / 31.9 / - 43.8 / 29.6 / - - / - / - 42.2 / 34.4 / - - / - / -
ASAL [43] 52.5 / 37.9 / - 44.9 / 32.1 / - 34.4 / - / - 39.2 / - / - - / - / -
TOT [35] 47.5 / 31.0 / - 40.6 / 30.0 / - 31.8 / - / - 47.4 / 42.8 / - 56.3 / 51.7 / -
TOT+ [35] 39.0 / 30.3 / - 45.3 / 32.9 / - 34.3 / - / - 44.5 / 48.2 / - 58.1 / 53.4 / -
UFSA (M) [71] - / - / - 43.2 / 30.5 / - - / - / - 47.8 / 34.8 / - - / - / -
UFSA (T) [71] 52.1 / 38.0 / - 49.6 / 32.4 / - 36.7 / 30.4 / - 55.8 / 50.3 / - 65.4 / 63.0 / -
ASOT [80] 56.1 / 38.3 / 18.6 52.9 / 35.1 / 24.7 46.2 / 37.4 / 24.9 59.3 / 53.6 / 30.1 70.4 / 68.0 / 45.9
HVQ [60] 54.4 / 39.7 / - 50.3 / 35.1 / - - / - / - - / - / - - / - / -
@@
Video Alignment Comparison Results. We now benchmark our VAOT and VASOT approaches against previous self-supervised video alignment methods and present quantitative results in Tab. 2. Firstly, it is evident from Tab. 2 that our VAOT approach achieves the best overall performance across all datasets, outperforming all competing methods, i.e., SAL [48], TCN [57], TCC [19], LAV [26], VAVA [45], and GTCC [17]. Especially, VAOT shows major improvements over VAVA [45] on the in-the-wild IKEA ASM dataset. The results confirm the advantage of our FGW formulation with a structural prior over the classical Kantorovich formulation with an optimality prior in VAVA [45]. Next, Fig. 4 shows some qualitative results, where VAOT retrieves all 5 correct frames with the same action (Hands at shoulder) as the query image, while LAV [26] obtains 3 incorrect frames (Hands above head), highlighted by red ovals. Lastly, similar to Fig. 3(d)(e), we find that action segmentation offers little benefit to video alignment in a multi-task learning setup, and our multi-task VASOT approach performs mostly similarly to our single-task VAOT approach in Tab. 2. This is likely because video alignment is a more complex problem involving finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation. Nevertheless, VASOT obtains mostly favorable results over prior works.
Action Segmentation Comparison Results. We test our VASOT approach against state-of-the-art self-supervised action segmentation methods and include quantitative results in Tab. 3. Firstly, from Tab. 3, VASOT consistently achieves the best results across all metrics and datasets, outperforming all competing methods, i.e., CTE [34], VTE [73], UDE [65], ASAL [43], TOT [35], UFSA [71], ASOT [80], and HVQ [60]. While our multi-task VASOT approach shows small gains over the single-task ASOT baseline [80] on Breakfast and YouTube Instructions, our improvements on 50 Salads and Desktop Assembly are substantial. The results validate the benefit of fusing video alignment with action segmentation and demonstrate that video alignment boosts action segmentation results notably in a multi-task learning setup. Moreover, Fig. 5 shows some qualitative results, where VASOT predicts segmentations which capture action boundaries more accurately and are more closely aligned with ground truth than ASOT [80].

### v3

@@ 3.1.3 Self-Supervised Learning
with $\tau>0$ denoting a temperature scaling parameter. In addition, pseudo-labels $\mathbf{T}^{\star}$ are obtained by solving the balanced FGW problem in Eq. 3 but with the augmented KOT cost matrix $\tilde{\mathbf{C}}=\mathbf{C}+\rho\mathbf{R}$, with $\rho\geq 0$. Here, we follow VAVA [liu2022learning] to impose the temporal prior $\mathbf{R}$, defined as $\mathbf{R}_{ij}=\left|i/N-j/M\right|$, which encourages the coupling $\mathbf{T}$ to have a banded diagonal shape and temporally nearby frames in $X$ to be matched with temporally adjacent frames in $Y$, yielding improved performance as seen in Sec. 4.1. Finally, our self-supervised learning loss is written as:
@@ 3.2 Joint Self-Supervised Video Alignment and Action Segmentation
Sec. 3.1 presents our self-supervised video alignment approach (VAOT) developed based on an FGW optimal transport formulation with a structural prior, which has previously been adopted for self-supervised action segmentation by ASOT [xu2024temporally]. Below we discuss differences between self-supervised video alignment and action segmentation, which lead to distinct design choices for VAOT and ASOT [xu2024temporally]. Firstly, as illustrated in Fig. 1, video alignment performs finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation, which causes our removal of $\mathbf{A}$ and our cost matrices $\{\mathbf{C},\mathbf{C}^{x},\mathbf{C}^{y}\}$ in Sec. 3.1 to be different from ASOT [xu2024temporally]. Secondly, video alignment generally has a more balanced assignment than action segmentation (the number of frames is much larger than the number of actions) and it is hard to balance multiple losses (the full unbalanced formulation adds two extra penalty terms). Thus, a balanced FGW formulation performs the best for VAOT, whereas a partial unbalanced FGW formulation is preferred in ASOT [xu2024temporally]. Thirdly, a background action class, to which background/redundant frames are assigned, is typically included in the $K$ actions for action segmentation, whereas it is not already defined for video alignment. Thus, VAOT adds a virtual frame to tackle background/redundant frames.
@@ 3.2.2 Self-Supervised Multi-Task Learning
Since both self-supervised video alignment and action segmentation exploit fine-grained temporal information in videos, we propose a self-supervised multi-task learning framework for joint video alignment and action segmentation. In particular, we combine our VAOT module for video alignment with ASOT [xu2024temporally] for action segmentation into a unified optimal transport-based approach (VASOT), which is illustrated in Fig. 2(b). Here, we update variable names in VAOT from $\mathbf{C},\mathbf{P},\mathbf{T}^{\star},\mathcal{L}$} to $\mathbf{C}^{xy},\mathbf{P}^{xy},\mathbf{T}^{xy\star},\mathcal{L}_{xy}$} respectively for video alignment between $X$ and $Y$, while introducing new variables $\mathbf{C}^{xa},\mathbf{P}^{xa},\mathbf{T}^{xa\star},\mathcal{L}_{xa}$} and $\mathbf{C}^{ya},\mathbf{P}^{ya},\mathbf{T}^{ya\star},\mathcal{L}_{ya}$} for action segmentation on $X$ and $Y$ respectively. The parameters $\bm{\theta}$ of the frame encoder and the action embeddings $\mathbf{A}$ are trained by using the below combination of self-supervised learning losses:
@@
where $w_{align}\geq 0$ and $w_{seg}\geq 0$ denote the weights for the video alignment loss $\mathcal{L}_{xy}$ and the action segmentation losses $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively, $\mathcal{L}_{xy}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xy}$ (computed between $\mathbf{X}$ and $\mathbf{Y}$) and pseudo-labels $\mathbf{T}^{xy\star}$ (obtained by VAOT), while $\mathcal{L}_{xa}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xa}$ (computed based on $\mathbf{X}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{xa\star}$ (derived by ASOT [xu2024temporally]) and $\mathcal{L}_{ya}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{ya}$ (computed between $\mathbf{Y}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{ya\star}$ (obtained by ASOT [xu2024temporally]). This is in contrast with VAOT or ASOT [xu2024temporally], where $\bm{\theta}$} is solely learned by using either $\mathcal{L}_{xy}$ or $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively. We find balancing $w_{align}=w_{seg}=1$ yields good results for both video alignment and action segmentation, as seen in Sec. 4.2. Our joint model requires training and storing a single model, saving both time and memory usage as compared to two single-task models. As we observe in Sec. 4.3, in a multi-task learning setting, action segmentation offers little benefit to video alignment performance, while video alignment boosts action segmentation results substantially.
@@ 4 Experiments
Datasets. We benchmark our VAOT and VASOT approaches for video alignment using three datasets, including monotonic datasets, i.e., Pouring [sermanet2018time] and Penn Action [zhang2013actemes], and in-the-wild dataset, i.e., IKEA ASM [ben2021ikea]. Pouring includes videos of humans pouring liquids and Penn Action comprises of videos of humans playing sports, while IKEA ASM videos capture humans assembling furniture. All methods have the same training and validation splits. Moreover, for action segmentation evaluation, we use four datasets, including in-the-wild datasets, i.e., Breakfast [kuehne2014language], 50 Salads [stein2013combining], and YouTube Instructions [alayrac2016unsupervised], and monotonic dataset, i.e., Desktop Assembly [kumar2022unsupervised]. Breakfast and 50 Salads videos show cooking activities, while YouTube Instructions consists of instructional videos and Desktop Assembly includes videos of an assembly activity. All methods are trained and tested on the same set of videos. For 50 Salads, we evaluate at two action granularity levels, i.e., Mid with 19 actions and Eval with 12 actions. Finally, for datasets with many activities, i.e., Penn Action, Breakfast, and YouTube Instructions, we train and test the methods per activity and report the average results.
@@
Implementation Details. For fair comparison purposes, our VAOT and VASOT approaches for video alignment utilize the same ResNet-50 encoder as recent self-supervised video alignment methods [dwibedi2019tcc, haresh2021learning, liu2022learning, donahue2024learning]. Similarly, our VASOT approach for action segmentation employs the same MLP encoder as state-of-the-art self-supervised action segmentation methods [kukleva2019unsupervised, kumar2022unsupervised, xu2024temporally]. Action embeddings $\mathbf{A}$ are initialized via $K$-Means, while the number of clusters $K$ is set to the ground truth value. We implement our methods in PyTorch [paszke2017automatic] and use ADAM optimization [kingma2014adam]. Please refer to our supplementary material for more details.
@@
Competing Methods. We compare our VAOT and VASOT approaches against prior self-supervised video alignment methods, namely SAL [misra2016shuffle], TCN [sermanet2018time], TCC [dwibedi2019tcc], LAV [haresh2021learning], VAVA [liu2022learning], and GTCC [donahue2024learning]. VAVA [liu2022learning], which integrates an optimality prior into a classical Kantorovich optimal transport, is the closest to our VAOT approach. Also, we test our VASOT approach against previous self-supervised action segmentation methods, namely CTE [kukleva2019unsupervised], VTE [vidalmata2021joint], UDE [swetha2021unsupervised], ASAL [li2021action], TOT [kumar2022unsupervised], UFSA [tran2024permutation], ASOT [xu2024temporally], and HVQ [spurio2024hierarchical]. ASOT [xu2024temporally] is the single-task baseline for action segmentation, which we adopt in Sec. 3.2 for our multi-task VASOT approach.
@@
Evaluation Metrics. To evaluate our VAOT and VASOT approaches for video alignment, we compute four metrics on the validation set, i.e., phase classification (Acc@{0.1,0.5,1.0}), phase progression (Progress), video alignment ($\tau$), and fine-grained frame retrieval (AP@{5,10,15}). Prior to that, we train the model on the training set and freeze it, and then train an SVM classifier or linear regressor on top of frozen features. Note that as mentioned in [dwibedi2019tcc, haresh2021learning, liu2022learning, donahue2024learning], Progress and $\tau$ are only defined for monotonic datasets and hence are not computed for IKEA ASM. Also, to test our VASOT approach for action segmentation, we calculate three metrics, i.e., mean over frames (MoF), F1 score (F1), and mean intersection over union (mIoU). Before that, we train the model, obtain predicted action segments, and perform Hungarian matching between predicted and ground truth action clusters.
@@ 4.2 Sensitivity Analysis Results
Here, we conduct sensitivity analyses on hyperparameters of VAOT and VASOT. Fig. 3 shows the results. We use Pouring in Figs. 3(a-e) and Desktop Assembly in Fig. 3(f). The results of $\epsilon$ are provided in our supplementary material.
Effect of $r$ and $\alpha$. From Fig. 3(a), Acc@1.0 remains stable and Progress shows small variation across all studied values of $r$, whereas $\tau$ is the most sensitive metric, peaking at $r=0.02$ and decreasing as $r$ increases. Similar observations can be made for $\alpha$ in Fig. 3(b), where Acc@1.0 and Progress are mostly stable, whereas $\tau$ fluctuates the most, performing the best with $\alpha=0.3$. Furthermore, we find that $r=0.02$ and $\alpha=0.3$ also work the best for the remaining datasets.
Effect of $\rho$ and $\zeta$. For $\rho$ in Fig. 3(c), it can be seen that Acc@1.0 and Progress remain mostly stable, whereas $\tau$ varies the most, performing the best at $\rho=0.35$. Similarly, for $\zeta$ in Fig. 3(d), we observe that $\tau$ is the most sensitive metric, peaking at $\zeta=0.5$, whereas Acc@1.0 and Progress remain steady across the analyzed value range of $\zeta$. Moreover, we notice that $\rho=0.35$ and $\zeta=0.5$ also yield the best results for the remaining datasets.
Effect of $w_{seg}$ and $w_{align}$. We study the relationship between action segmentation and video alignment in our multi-task VASOT approach by varying their weights $w_{seg}$ and $w_{align}$ in Eq. 7. In particular, we report alignment results with various $w_{seg}$ values (while keeping $w_{align}=1$) in Fig. 3(e), and segmentation results with various $w_{align}$ values (while keeping $w_{seg}=1$) in Fig. 3(f). It can be seen from Fig. 3(e) that alignment results drop as $w_{seg}$ increases and the best overall alignment performance is obtained with $w_{seg}=1$. In contrast, it is clear from Fig. 3(f) that segmentation results improve as $w_{align}$ increases and become steady after $w_{align}$ reaches $1$. Therefore, in a multi-task learning setup, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance notably. Moreover, with balancing $w_{seg}=w_{align}=1$, VASOT achieves good results for both video alignment and action segmentation. For the next section, we set $w_{seg}=w_{align}=1$ for VASOT.
@@ 4.3 State-of-the-Art Comparison Results
Pouring SAL [misra2016shuffle] 85.68 87.84 88.02 74.51 73.31 84.05 83.77 83.79
TCN [sermanet2018time] 89.19 90.39 90.35 80.57 86.69 83.56 83.31 83.01
TCC [dwibedi2019tcc] 89.23 91.43 91.82 80.30 85.16 87.16 86.68 86.54
LAV [haresh2021learning] 91.61 92.82 92.84 80.54 85.61 89.13 89.13 89.22
VAVA [liu2022learning] 91.65 91.79 92.45 83.61 87.55 90.05 89.92 90.17
GTCC [donahue2024learning] 71.20 89.20 93.50 85.80 88.10 - - -
@@
Penn Action SAL [misra2016shuffle] 74.87 78.26 79.96 59.43 63.36 76.04 75.77 75.61
TCN [sermanet2018time] 81.99 83.67 84.04 67.62 73.28 77.84 77.51 77.28
TCC [dwibedi2019tcc] 81.26 83.35 84.45 67.26 73.53 76.74 76.27 75.88
LAV [haresh2021learning] 83.56 83.95 84.25 66.13 80.47 79.13 78.98 78.90
VAVA [liu2022learning] 83.89 84.23 84.48 70.91 80.53 81.52 80.47 80.67
GTCC [donahue2024learning] 78.30 81.20 81.30 70.80 88.30 - - -
@@
IKEA ASM SAL [misra2016shuffle] 22.94 23.43 25.46 - - 14.28 14.04 14.10
TCN [sermanet2018time] 22.51 25.47 25.88 - - 17.37 17.03 16.96
TCC [dwibedi2019tcc] 22.70 25.04 25.63 - - 18.03 17.53 17.20
LAV [haresh2021learning] 23.19 25.47 25.54 - - 20.14 19.35 19.21
VAVA [liu2022learning] 29.12 29.95 29.10 - - 26.42 25.73 25.80
@@
Full-Dataset Evaluation CTE* [kukleva2019unsupervised] 41.8 / 26.4 / - 39.0 / 28.3 / - 30.2 / - / - 35.5 / - / - 47.6 / 44.9 / -
CTE† [kukleva2019unsupervised] 47.2 / 27.0 / 14.9 35.9 / 28.0 / 9.9 30.1 / 25.5 / 17.9 35.0 / 35.5 / 21.6 - / - / -
VTE [vidalmata2021joint] 48.1 / - / - - / 29.9 / - 24.2 / - / - 30.6 / - / - - / - / -
UDE [swetha2021unsupervised] 47.4 / 31.9 / - 43.8 / 29.6 / - - / - / - 42.2 / 34.4 / - - / - / -
ASAL [li2021action] 52.5 / 37.9 / - 44.9 / 32.1 / - 34.4 / - / - 39.2 / - / - - / - / -
TOT [kumar2022unsupervised] 47.5 / 31.0 / - 40.6 / 30.0 / - 31.8 / - / - 47.4 / 42.8 / - 56.3 / 51.7 / -
TOT+ [kumar2022unsupervised] 39.0 / 30.3 / - 45.3 / 32.9 / - 34.3 / - / - 44.5 / 48.2 / - 58.1 / 53.4 / -
UFSA (M) [tran2024permutation] - / - / - 43.2 / 30.5 / - - / - / - 47.8 / 34.8 / - - / - / -
UFSA (T) [tran2024permutation] 52.1 / 38.0 / - 49.6 / 32.4 / - 36.7 / 30.4 / - 55.8 / 50.3 / - 65.4 / 63.0 / -
ASOT [xu2024temporally] 56.1 / 38.3 / 18.6 52.9 / 35.1 / 24.7 46.2 / 37.4 / 24.9 59.3 / 53.6 / 30.1 70.4 / 68.0 / 45.9
HVQ [spurio2024hierarchical] 54.4 / 39.7 / - 50.3 / 35.1 / - - / - / - - / - / - - / - / -
@@
Video Alignment Comparison Results. We now benchmark our VAOT and VASOT approaches against previous self-supervised video alignment methods and present quantitative results in Tab. 2. Firstly, it is evident from Tab. 2 that our VAOT approach achieves the best overall performance across all datasets, outperforming all competing methods, i.e., SAL [misra2016shuffle], TCN [sermanet2018time], TCC [dwibedi2019tcc], LAV [haresh2021learning], VAVA [liu2022learning], and GTCC [donahue2024learning]. Especially, VAOT shows major improvements over VAVA [liu2022learning] on the in-the-wild IKEA ASM dataset. The results confirm the advantage of our FGW formulation with a structural prior over the classical Kantorovich formulation with an optimality prior in VAVA [liu2022learning]. Next, Fig. 4 shows some qualitative results, where VAOT retrieves all 5 correct frames with the same action (Hands at shoulder) as the query image, while LAV [haresh2021learning] obtains 3 incorrect frames (Hands above head), highlighted by red ovals. Lastly, similar to Fig. 3(e), we find that action segmentation offers little benefit to video alignment in a multi-task learning setup, and our multi-task VASOT approach performs mostly similarly to our single-task VAOT approach in Tab. 2. This is likely because video alignment is a more complex problem involving finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation. Nevertheless, VASOT obtains mostly favorable results over prior works.
Action Segmentation Comparison Results. We test our VASOT approach against state-of-the-art self-supervised action segmentation methods and include quantitative results in Tab. 3. Firstly, from Tab. 3, VASOT consistently achieves the best results across all metrics and datasets, outperforming all competing methods, i.e., CTE [kukleva2019unsupervised], VTE [vidalmata2021joint], UDE [swetha2021unsupervised], ASAL [li2021action], TOT [kumar2022unsupervised], UFSA [tran2024permutation], ASOT [xu2024temporally], and HVQ [spurio2024hierarchical]. While our multi-task VASOT approach shows small gains over the single-task ASOT baseline [xu2024temporally] on Breakfast and YouTube Instructions, our improvements on 50 Salads and Desktop Assembly are substantial. The results validate the benefit of fusing video alignment with action segmentation and demonstrate that video alignment boosts action segmentation results notably in a multi-task learning setup. Moreover, Fig. 5 shows some qualitative results, where VASOT predicts segmentations which capture action boundaries more accurately and are more closely aligned with ground truth than ASOT [xu2024temporally].